# PAE MobiLLM: Privacy-Aware and Efficient LLM Fine-Tuning on the Mobile Device via Additive Side-Tuning

### v1

@@ 6.4. Advantages of LLM FT Acceleration
  RoBERTa-large LoRA-L - 2.66 - 10.07 - 2.66 10.07
BitFit-L - 2.54 - 9.88 - 2.54 9.88
MobiLLM-L - 1.73 - 7.37 - 1.73 7.37
Full-FT-SL 2.22 0.30 4.44 1.25 0.06 2.58 5.75
LoRA-SL 2.22 0.21 4.44 0.58 0.06 2.49 5.08
BitFit-SL 2.22 0.19 4.44 0.53 0.04 2.45 5.01
MobiLLM 13.36 1.07 26.72 4.63 0.01 13.36 26.72
PAE MobiLLM 0.05 1.09 0.10 4.61 0.01 1.09 4.61
Impact of Wireless Transmission Speed. Figure 10 illustrates the per-batch execution time across methods under different wireless transmission speeds. Since the total execution time in BitFit-SL and MobiLLM is predominantly affected by the communication time, they suffer severe performance degradation under low-speed transmission conditions. For example, BitFit-SL shows an 8.75×8.75$\times$8.75 × slowdown at 10Mbps vs. 100Mbps. In contrast, PAE MobiLLM maintains stable total latency across different transmission speeds, demonstrating almost no variations from 10 Mbps to 100 Mbps. Such resilience stems from two facts: 1) single-token transmission takes very short time and thus minimizes the impact of wireless transmission speeds, and 2) parallelized scheduling overlaps communication phases with device-side computation, hiding transmission delays with on-device computation time that remains unaffected by transmission speed fluctuations.
7. Discussions and Future Work
The proposed PAE MobiLLM design demonstrates significant advancements in mobile LLM FT, but several practical considerations may merit deeper exploration to broaden its applicability.
Accuracy and Efficiency Trade-off: While PAE MobiLLM significantly improves the time, communication, and computational efficiency of on-device LLM fine-tuning, its lightweight side-network architecture and selective activation transmission strategy yield minor accuracy trade-offs (e.g., 0.4% and 1.2% reductions on QNLI tasks compared to SST2 and MobiLLM methods respectively). Our future work will explore adaptive mechanisms to balance efficiency and accuracy for precision-sensitive scenarios.
Server Storage: The server-side storage demands for caching activation data pose scalability challenges in resource constrained cloud environments. Techniques such as activation-aware quantization (e.g., mixed 2-4 bit precision per layer) and intelligent lifecycle management—where stale or low-impact activation pairs are automatically purged—could align storage costs with operational budgets without compromising adaptation quality.
Optimizing Side-Tuning: The server-side training pipeline also holds untapped optimization potential. Inspired by large-batch training strategies, dynamically scaling batch dimensions based on real-time GPU memory utilization could accelerate convergence while maintaining stability. For instance, progressively increasing batch sizes from 32 to 1024 following linear scaling rules may exploit hardware parallelism more effectively.
Mobile AI System Heterogeneity: Lastly, the growing heterogeneity of mobile hardware (e.g., varying CPU/TPU capabilities across devices) necessitates platform-agnostic computation strategies. Developing energy-proportional training protocols — where layer-wise computations adapt to device-specific power budgets via dynamic frequency scaling — could democratize efficient fine-tuning across diverse edge ecosystems. These directions collectively aim to transform theoretical advancements into deployable solutions for the evolving edge intelligence landscape.
8. Conclusion
@@
This paper has presented PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via additive side-tuning. By decoupling the frozen backbone model on the mobile device from server-side side-network tuning, PAE MobiLLM offloads computation and memory-intensive workloads to the server while enabling practical LLM FT on mobile devices. To improve computing efficiency, PAE MobiLLM integrates activation caching on the server side to remove iterative on-device computations by reusing historical activations. To improve communication efficiency, PAE MobiLLM exploits one-token activation selection to compress activations and reduce transmission workloads. Aware of label and fine-tuned model privacy, PAE MobiLLM introduces the additive adapter side-network that isolates label-sensitive operations from the server. Experimental results show that PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × acceleration in LLM FT convergence, while keeping data, training labels, and fine-tuned models on the mobile device. This work provides a deployable LLM FT solution to the resource-constrained mobile AI device whose computation/communication efficiency and data privacy are essential.

### v2

@@ Introduction
Table 3: Per-Batch Execution Time Breakdown (task:SST-2@OPT-1.3B).
Per-Batch Transmission Efficiency. Per-batch transmission measurements for OPT-1.3B (Table 1) reveal fundamental architectural differences. Split learning based methods necessitate dual activation-gradient exchanges and result in 33.4MB/batch communication load, while MobiLLM’s further escalates this burden due to its full-layer activation transmissions. In contrast, PAE MobiLLM leverages the token selector to identify and transmit only the activation dimensions involved in the loss computation. Compared to split learning, it reduces per-batch traffic by at least 27.8% in generation task and even up to 97% in classification tasks, which makes such device-server collaboration based LLM FT feasible in low-speed or unstable wireless networks.
Total Communication Load. As demonstrated in the Table 1, all the other methods require iterative interactions with the server throughout the LLM FT process and lead to accumulated communication load proportional to the number of epoch. By contrast, PAE MobiLLM processes the entire training dataset within the first epoch. It has no need for subsequent activation transmissions and only has a fixed initial cost. This promisingly enables precise estimation of total bandwidth requirements, which is critical for distributed training orchestration.
Advantages of LLM FT Acceleration
We evaluate the training efficiency of PAE MobiLLM by measuring the total time required to reach target accuracy across diverse configurations involving two model architectures (OPT, RoBERTa), two tasks (MRPC, QNLI), and two mobile devices (NVIDIA Xavier, CPU-only laptop), as shown in Fig. 5. For clarity, we simplify the presentation by retaining only BitFit’s training curves (due to its similarity to LoRA) and excluding memory-intensive local FT baselines that exceed the mobile device’s memory capacity.
PAE MobiLLM is much faster than conventional split learning and server-assisted LLM FT baselines. Taking the FT of the MRPC task with the OPT-1.3B model (on Xavier) as an example, PAE MobiLLM is $15.5\times$ faster than the second-fastest approach, BitFit-SL. Particularly, PAE MobiLLM demonstrates a two-phase acceleration pattern throughout the FT process. In the initial phase, PAE MobiLLM’s parallel scheduling design allows the mobile device to process subsequent data batches while the server simultaneously trains side networks using current-batch activations. This eliminates the idle time of the mobile device/server during activation transmissions and server computations, which is a delay bottleneck in traditional split learning frameworks. Besides, using token selector before transmission in PAE MobiLLM remarkably reduces the transmission delay, which further speeds up the training. After the first epoch, PAE MobiLLM leverages server-side caching to bypass redundant device-side computation and communication for identical sample activations. By leveraging the server’s computational resources to perform independent iterative training, PAE MobiLLM achieves faster LLM FT while maintaining stable convergence, which outperforms iterative recomputation and communication based LLM FT approaches.
Breakdown of Per-Batch Execution Time. Table 3 compares the time costs, including communication time $t_{1}$, on-device computation time $t_{2}$, and server-side processing time $t_{3}$, for forward and backward passes across different approaches. Split learning based methods execute all the stages sequentially and result in cumulative latency of the three parts. While MobiLLM supports device-server parallel computation, its heavy activation transmission burden dominates its total LLM FT latency, especially when processing long sequences or large batches. Unlike those designs, PAE MobiLLM alleviates this bottleneck through token selection, reduces per-batch data transmission and makes transmission latency negligible. For the on-device computing time, both MobiLLM and PAE MobiLLM remove backward propagation workloads from the mobile devices, which only need to conduct forward inference computations. That helps to simplify on-device operations, making MobiLLM and PAE MobiLLM naturally compatible with emerging inference accelerators tailored for mobile devices (e.g., mobile TPUs). As shown in Table 3, PAE MobiLLM consistently achieves the shortest per-batch execution time across all tested devices and model scales, demonstrating superior training efficiency.
LLM FT Performance Analysis
PAE MobiLLM introduces three key components beyond the MobiLLM framework, i.e., the additive side network, the token selector and the server cache, respectively. We conducted ablation studies to assess the contribution of each component to LLM FT performance. For the GLUE benchmark, we report the average score, which combines accuracy on MNLI, QQP, QNLI, SST-2, MRPC, and RTE, Pearson correlation on STS-B, and Matthews correlation on CoLA. For the DialogSum dataset, we use ROUGE scores (R1/R2/RL) as the accuracy metric. As shown in Table 4, both the server cache and the token selector significantly improve training efficiency without degrading LLM FT performance. The additive side network also consistently maintains competitive performance across different datasets. The experimental results demonstrate that the proposed PAE MobiLLM achieves excellent training efficiency and enhanced data privacy without sacrificing LLM FT performance.
Model OPT-1.3B RoBERTa-Large
Data
GLUE
Dialogsum
(R1/R2/RL)
GLUE
PAE MobiLLM 81.1 21.1 / 7.9 / 17.1 86.8
w/o. ASN 81.4 21.3 / 7.5 / 17.6 86.9
w/o. TS 81.2 21.1 / 7.7 / 17.0 86.7
w/o. SC 81.1 21.0 / 7.7 / 17.1 86.8
w/o. ASN/ TS/ SC 81.4 21.3 / 7.6 / 17.5 86.6
Table 4: Ablation studies of the proposed PAE MobiLLM designs (ASN: Additive Side Network, TS: Token Selector, SC: Sever Cache, w/o. : Without).
Conclusion
@@
This paper has presented PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via additive side-tuning. By decoupling the frozen backbone model on the mobile device from server-side side-network tuning, PAE MobiLLM offloads computation and memory-intensive workloads to the server while enabling practical LLM FT on mobile devices. To improve computing efficiency, PAE MobiLLM integrates activation caching on the server side to remove iterative on-device computations by reusing historical activations. To improve communication efficiency, PAE MobiLLM exploits token selection to compress activations and reduce transmission workloads. Aware of label privacy, PAE MobiLLM introduces the additive adapter side-network that isolates label-sensitive operations from the server. Experimental results show that PAE MobiLLM achieves at least a $1.79\times$ reduction in device-side computation, a $17.06\times$ reduction in communication cost, and a $5.25\times$ acceleration in LLM FT convergence, while keeping data and labels private.